# A Prompt-Engineered Large Language Model, Deep Learning Workflow for Materials Classification

### v1

@@ A Prompt-Engineered Large Language Model, Deep Learning Workflow
Siyu Liu et al.
@@ 1.1 Glossary
Attention: a mechanism enabling models to focus on specific parts of the input sequence during processing, allowing them to weigh the significance of different words when generating output. This attention mechanism assists in capturing long-range dependencies and enhancing performance in various natural language processing (NLP) tasks [52].
@@
Feature Engineering: the process of selecting, transforming, and creating new features from raw data to improve machine leanring (ML) model performance. It involves identification of relevant and informative input variables and their preparation for accurate predictions or classifications [57].
Feed Forward: refers to a deep learning framework, also known as a multi-layer perceptron (MLP), that defines a mapping , and learns the value of parameter to achieve the best function approximation [58].
Few-shot Learning: a ML approach where models are trained to make accurate predictions with limited examples per class [59].
Gemini: Google’s multimodal large language model (LLM), the first to outperform human experts on MMLU (Massive Multitask Language Understanding), a popular method for testing AI models’ knowledge and problem-solving abilities.
Gradient Boosting Decision Tree (GBDT): also known as Gradient Boosting Machines (GBM), is a popular ensemble learning method combining decision trees with gradient boosting. GBDT continuously fits new models for more accurate estimates of response variables. The principle behind this algorithm is to construct new base learners that maximizes correlation with the negative gradient of the ensemble’s loss function [60].
@@
In-context Learning: refers to enhancing LLM performance using a few examples provided in the input context [62].
@@
Large Language Model: large language model (LLM) is a kind of DL model trained on large amounts of text data to learn statistical patterns of natural language [64].
@@
Pre-trained Model: a model trained on a large corpus of data and can be fine-tuned to solve various tasks [67].
@@
Self-instruct Learning: a method for improving the instruction-following capabilities of pre-trained language models by bootstrapping off their own generations [68].
Supervised Classification: a ML task where the goal is to categorize input data into predefined classes or categories based on labeled training examples [69].
@@
Transfer Learning: a ML technique employing a pre-trained model as the starting point for a new related task, instead of training a model from scratch. It allows for different domains, tasks, and distributions in training and testing [72].
@@ 1.2.1 Data processing of metallic glasses dataset
As mentioned in the “Methods” section, our initial dataset consists of 5,577 samples. To train different BERT models, it is necessary to divide it into a training and a test set. We employed a stratified sampling method for this purpose, ensuring a representative and balanced representation of the different subgroups [73], which enhances the quality and reliability of sample. Based on the overall 80:20 split ratio for the training and test sets, we have performed non-repetitive stratified sampling on three alloy data categories. In the end, we obtained a training set with 4,460 samples and a test set with 1,117 samples, maintaining the 80:20 ratio for each category.
It should be noted that this section aims to establish the relationship between alloy composition and labels. The textual data generated through Gemini-pro is used to train different BERT classification models. The data obtained in this section is stored in the “original data” folder, with the data format shown in Table S1.
@@ 1.2.2 Prompt design and textual data generation
Upon acquiring the alloy composition data, we proceeded with prompt design. Prompt engineering is a prominent direction in the field of LLMs and is considered a method that significantly improve data generation effectiveness for LLMs [74]. Techniques such as few-shot learning [53], CoT [53], synergizing reasoning and acting (ReAct) [75], and retrieval-augmented generation (RAG) [76] are deemed effective prompt engineering methods. Due to the additional knowledge sources required for the latter two methods, we tested direct inquiry, few-shot, and CoT, and compared them with our MetalPrompt. Figs. S8, S9, and S10 are templates for three benchmark methods, while Fig. S4 shows the comparison results. Our “MetalPrompt” achieved the best performance.
For the prompt inquiry setting with the LLM, we used Top-K = 1 and Temperature = 0 as default parameters. As shown in Fig. S11 and Fig. S12, a smaller Top-K and lower temperature indicate more reliable model output. Top-K = 1 implies that the model will perform greedy decoding and select the most probable value. Unless specifically explained, the parameters of the language model mentioned below are consistent with these settings.
Using the above parameter settings and “MetalPrompt” as the input template, we generated textual data for 5,577 alloy compositions. The replacement of alloy elements in the template was assisted by the langchain library. Simultaneously, based on the alloy composition category relationship of the original training and test dataset, we labeled the generated textual data with the same category. This dataset is used for training the classification model below, with the partitioning of training and testing sets consistent with the original data.
@@ 1.2.3 MgBERT training and evaluation
BERT model and attention mechanism: BERT, introduced by Google researchers in 2018 [27], is a pre-training language representation method that significantly impacted the field of NLP. Traditional language models analyzed text data unidirectionally, either from left to right or right to left, limiting their understanding of language context. BERT, employing the Transformer architecture, processes each word concerning all other words in a sentence, rather than sequentially. Owing to its state-of-the-art results in numerous tasks, BERT was selected as the foundation for our classification model.
@@
Fig. 4c in the main text shows the basic BERT architecture, comprising data pre-processing, input encoding, model training, post-processing, and output results. During data processing, tokenizers divide sentences or words into individual tokens, such as marking “for example” as “for” and “example”. Then, as demonstrated in Fig. S13, the encoding method transforms pre-processed text data into the input representation of the model. For instance, “for” is converted to a vector . Model training is crucial to BERT, as it uses multi-head attention to extract essential features from input embeddings. Attention mechanisms enhance model performance in NLP and other sequence data processing tasks by focusing on the most relevant parts of the input sequence. Models can learn to assign various attention weights to input information from different positions, concentrating on crucial aspects when processing input sequences. This mechanism allows the model to capture long-distance dependencies and important patterns in sequences more effectively, thus improving its performance in processing sequence data [77]. Fig. S14 shows a schematic diagram of the operation of the attention mechanism. Multi-head attention, compared to a single module, employs multiple attention modules in the same calculation, extracting different feature information. The calculation formula for attention score is as follows:
Where .
@@
In Equation S1, refers to the queries and dimension keys. , , and refer to the input query, key, and value, respectively. refers to the softmax function.
Pre-trained models selection: We compared three pre-trained models based on BERT, including the basic BERT model [27], Longformer supporting longer input text lengths [78], and MatSciBERT trained on materials science texts [42].
@@
All models employ default parameter settings; however, to expedite calculations, the maximum input length for Longformer is limited to 1200 tokens, while other models have a default maximum length of 512. To ensure reproducibility, the random seed is fixed at 42. All models are trained using cross-entropy loss and optimized with the Adam optimizer for parameters. All training was conducted on a 32G Nvidia V100. Table S2 presents the basic hyperparameters for model training. Due to its larger embedding layer requiring more graphics memory, Longformer has a slightly smaller batch size.
@@
MgBERT embedding and training: Upon the training of MgBERT, we found that the file with the longest tokens is ’Fe68.3C6.9Si2.5B6.7P8.8Cr2.2Mo2.5Al2.1.txt’ with 859 tokens, and 300 files have more than 512 tokens. Our basic model, MatSciBERT, only supports inputs with a maximum token length of 512. Considering the potential inconsistency in text input length for different materials and scalability requirements, we designed a variable-length MgBERT position embedding layer that does not necessitate complete retraining. The following Algorithm 1 is a pseudocode representation of its principle.
@@
For the comparison of three embedding methods, we ran 16 epochs with a learning rate of 3e-5 and a batch size of 26. In Figs. 4g, h of the main text, only the first 15 epochs are displayed due to the loss output retaining the first three decimal places. In the 16th epoch, some models exhibit training set loss 0.001. The complete training information can be found in the “train log” folder.
@@
For the final MgBERT model, we continued training with the highest classification accuracy model weight of Method 3 at a learning rate of 2e-5 and compared it to the case with 1e-5. The highest accuracy of 88.5% was achieved when the learning rate was 2e-5 and the epoch was 4. We utilized the model weights trained under these parameters as the final MgBERT model.
@@
Model evaluation: During the model evaluation stage, we compared some ML models, with their concepts outlined in the “Glossary” section. Model implementation is based on the “scikit-learn” software and employs its default hyperparameters for training. The random seed is fixed at 42 for reproducibility. The training and test set rely on stratified sampling, as described in the “Methods” section of the main text.
@@ 1.2.4 Model interpretability
We used the LIME model [44] for model interpretation. LIME utilizes a trained local surrogate model to explain individual predictions. For a black-box model of interest, LIME samples around the instance of interest and collects perturbed instances along with their black-box model predictions. A local interpretable model (e.g. linear regression, decision tree) is then trained on this new dataset to approximate the behavior of the black-box model locally. Fig. S15 from [44] shows its working principle, wherein the dashed line is a trained linear model used to explain the black-box model with complex boundaries in the red and blue part of the graph. The hyperparameters we set here is to perform 1,000 samples and filter the top 50 contributing features. Considering the maximum token length does not exceed 900, this choice is reasonable under the limitation of computational rate, as the training time increases exponentially with the increase of sampling times.
Additionally, we visualized the attention score of the [CLS] (classify) token in the last layer. Different attention scores represent the degree of importance that a specific token places on other tokens, with larger attention scores indicating higher importance. The classifier processes the value of the last layer’s [CLS] token, compressing this 768-dimensional vector into three dimensions. Each dimension represents the probability of different categories. Therefore, calculating the focus on the previous layer assists in interpreting the model.
@@ 3 Supplementary Tables
glass forming category |
composition |
Ribbon |
Ag20Al25La55 |
BMG |
Cu55Zr42.5Ga2.5 |
NR |
Ag6Ce8Cu86 |
model name |
hyperparameters |
BERT |
Epoch: 8; Learning rate: 3e-5; Batch size: 26 |
longformer |
Epoch: 8; Learning rate: 3e-5; Batch size: 12 |
MatSciBERT |
Epoch: 8; Learning rate: 3e-5; Batch size: 26 |

### v2

@@ A Prompt-Engineered Large Language Model, Deep Learning Workflow
Siyu Liu, Tongqi Wen, A. S. L. Subrahmanyam Pattamatta, David J. Srolovitz
@@ 1.1 Glossary
Attention: a mechanism enabling models to focus on specific parts of the input sequence during processing, allowing them to weigh the significance of different words when generating output. The attention mechanism assists in capturing long-range dependencies and enhancing performance in various natural language processing (NLP) tasks [57].
@@
Feature Engineering: selecting, transforming, and creating new features from raw data to improve machine learning (ML) model performance. It involves identifying relevant and informative input variables and preparing them for accurate predictions or classifications [62].
Feed Forward: refers to a deep learning framework, also known as a multi-layer perceptron (MLP), that defines a mapping and learns the value of parameter to achieve the best function approximation [63].
Few-shot Learning: an ML approach where models are trained to make accurate predictions with limited examples per class [64].
Gemini: Google’s multimodal LLM, the first to outperform human experts on MMLU (Massive Multitask Language Understanding), a popular method for testing the knowledge and problem-solving abilities of AI models [53].
Gradient Boosting Decision Tree (GBDT): also known as Gradient Boosting Machines (GBM), is a popular ensemble learning method combining decision trees with gradient boosting. GBDT continuously fits new models for more accurate estimates of response variables. The principle behind this algorithm is to construct new base learners that maximize correlation with the negative gradient of the ensemble’s loss function [65].
@@
In-context Learning: refers to enhancing LLM performance using a few examples in the input context [67].
@@
Large Language Model: large language model (LLM) is a class of DL models trained on large amounts of text data to learn statistical patterns of natural language [69].
@@
Pre-trained Model: a model trained on a large corpus of data that may be fine-tuned to solve various tasks [72].
@@
Self-Instruct Learning: a method for improving the instruction-following capabilities of pre-trained language models by bootstrapping off their generations [73].
Supervised Classification: an ML task that categorizes input data into predefined classes or categories based on labeled training examples [74].
@@
Transfer Learning: an ML technique employing a pre-trained model as the starting point for a new related task instead of training a model from scratch. It allows for different domains, tasks, and distributions in training and testing [77].
@@ 1.2.1 Data processing of metallic glasses dataset
As mentioned in the “Methods” section, our initial dataset consists of 5,577 samples. To train different BERT models, it is necessary to divide it into a training and a test set. We employed a stratified sampling method for this purpose, ensuring a representative and balanced representation of the different subgroups [78], which enhances the quality and reliability of the sampling. We performed non-repetitive stratified sampling on three alloy data categories based on the overall 80:20 split ratio for the training and test sets. Ultimately, we obtained a training set with 4,460 samples and a test set with 1,117 samples, maintaining the 80:20 ratio for each category.
The textual data generated through Gemini-pro is used to train different BERT classification models. The data format of obtained data is shown in Table S1.
@@ 1.2.2 Prompt design and textual data generation
Upon acquiring the alloy composition data, we proceeded with prompt design. Prompt engineering is a prominent direction in the field of LLMs and is considered a method that significantly improves data generation effectiveness for LLMs [79]. Techniques such as few-shot learning [58], CoT [58], synergizing reasoning and acting (ReAct) [80], and retrieval-augmented generation (RAG) [81] are effective prompt engineering methods. Due to the additional knowledge sources required for the latter two methods, we tested direct inquiry, few-shot, and CoT, and compared them with our MetalPrompt. Figs. S8, S9, and S10 are templates for three benchmark methods, while Fig. S4 shows the comparison results. Our “MetalPrompt” achieved the best performance.
For the prompt inquiry setting with the LLM, we used Top-K = 1 and Temperature = 0 as default parameters. As shown in Figs. S11 and S12, a smaller Top-K and lower Temperature indicate more reliable model output. Top-K = 1 implies that the model performs greedy decoding and selects the most probable value. Unless specifically explained, the language model parameters mentioned below are consistent with these settings.
Using the above parameter settings and “MetalPrompt” as the input template, we generated textual data for 5,577 alloy compositions. The input of alloy elements in the template was assisted by the langchain [50] library. Simultaneously, based on the alloy composition category relationship of the original training and test dataset, we labeled the generated textual data with the same category. This dataset is used for training the classification model below, partitioning training, and testing sets consistent with the original data.
@@ 1.2.3 MgBERT training and evaluation
BERT model and attention mechanism: BERT [27] is a pre-training language representation method that has had significantly impact in the NLP field. Traditional language models analyzed text data unidirectionally, either from left to right or right to left, limiting their understanding of language context. Using the Transformer architecture, BERT processes each word concerning all other words in a sentence rather than sequentially. BERT was selected as the foundation for our classification model due to its state-of-the-art results in numerous tasks.
@@
Fig. 4c in the main text shows the basic BERT architecture, comprising data pre-processing, input encoding, model training, post-processing, and output results. During data processing, tokenizers divide sentences or words into individual tokens, such as replacing “for example” with “for” and “example”. Then, as demonstrated in Fig. S13, the encoding method transforms pre-processed text data into the input representation of the model. For instance, “for” is converted to a vector . Model training is crucial to BERT and we use multi-head attention to extract essential features from input embeddings. Attention mechanisms enhance model performance in NLP and other sequence data processing tasks by focusing on the most relevant parts of the input sequence. Models can learn to assign various attention weights to input information from different positions, concentrating on crucial aspects when processing input sequences. This mechanism allows the model to capture long-distance dependencies and important patterns in sequences more effectively, thus improving its performance in processing sequence data [82]. Fig. S14 shows a schematic diagram of the operation of the attention mechanism. Compared to a single module, multi-head attention employs multiple attention modules in the same calculation, extracting different feature information. The calculation formula for attention score is as follows:
@@
where , refers to the queries and dimension keys, softmax refers to the softmax function, and , , and refer to the input query, key, and value, respectively.
@@
Pre-trained models selection: We compared three pre-trained models based on BERT, including the basic BERT model [27], Longformer supporting longer input text lengths [83], and MatSciBERT trained on materials science texts [42]. All models employ default parameter settings; however, to expedite calculations, the maximum input length for Longformer is limited to 1200 tokens, while other models have a default maximum length of 512. To ensure reproducibility, the random seed is fixed at 42. All models are trained using cross-entropy loss and optimized with the Adam optimizer for parameters. All training was conducted on a 32G Nvidia V100. Table S2 presents the basic hyperparameters for model training. Due to its larger embedding layer requiring more graphics memory, Longformer has a slightly smaller batch size.
@@
MgBERT embedding and training: In training of MgBERT we found that the file with the longest tokens is “Fe68.3C6.9Si2.5B6.7P8.8Cr2.2Mo2.5Al2.1.txt” with 859 tokens, and 300 files have more than 512 tokens. Our basic model, MatSciBERT, only supports inputs with a maximum token length of 512. Considering the potential inconsistency in text input length for different materials and scalability requirements, we designed a variable-length MgBERT position embedding layer that does not necessitate complete retraining. The following Algorithm 1 is a pseudocode representation of its principle.
@@
To compare three embedding methods, we ran 16 epochs with a learning rate and a batch size 26. In Figs. 4g, h of the main text, only the first 15 epochs are displayed due to the loss output retaining the first three decimal places. In the epoch, some models exhibit training set loss 0.001.
@@
For the final MgBERT model, we continued training with the highest classification accuracy model weight of Method 3 at a learning rate of and compared it to the case with . The highest accuracy of 88.5% was achieved when the learning rate was , and the epoch was 4. We utilized the model weights trained under these parameters as the final MgBERT model.
@@
Model evaluation: During the model evaluation stage, we compared three ML models – their concepts are outlined in the “Glossary”. The model implementation is based on the “scikit-learn” software and employs its default hyperparameters for training. The random seed is fixed at 42 for reproducibility. The training and test set rely on stratified sampling, as described in the “Methods” section of the main text.
@@ 1.2.4 Model interpretability
We used LIME [44] for model interpretation. LIME utilizes a trained local surrogate model to explain individual predictions. For a black-box model of interest, LIME samples around the instance of interest and collects perturbed instances and their black-box model predictions. A local interpretable model (e.g., linear regression, decision tree) is then trained on this new dataset to approximate the behavior of the black-box model locally. Fig. S15 from [44] shows its working principle, wherein the dashed line is a trained linear model used to explain the black-box model with complex boundaries in the red and blue part of the graph. We set hyperparameters to sample 1,000 cases and filter the top 50 contributing features. Considering the maximum token length does not exceed 900, this choice is reasonable under the limitation of computational rate, as the training time increases exponentially with increasing sampling times.
Additionally, we visualized the attention score of the [CLS] (classify) token in the last layer. Different attention scores represent the degree of importance a specific token places on other tokens, with larger attention scores indicating higher importance. The classifier processes the value of the last layer [CLS] token, compressing this 768-dimensional vector into three dimensions; each dimension represents the probability of different categories. Therefore, calculating the focus on the previous layer assists in interpreting the model.
@@ 3 Supplementary Tables
| glass forming category | composition |
| Ribbon | Ag20Al25La55 |
| BMG | Cu55Zr42.5Ga2.5 |
| NR | Ag6Ce8Cu86 |
| model name | hyperparameters |
| BERT | Epoch: 8; Learning rate: 3e-5; Batch size: 26 |
| Longformer | Epoch: 8; Learning rate: 3e-5; Batch size: 12 |
| MatSciBERT | Epoch: 8; Learning rate: 3e-5; Batch size: 26 |